40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 23 13:30:06 2012, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
    Reply  Mon Jul 23 18:01:30 2012, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
    Reply  Thu Mar 14 16:41:48 2013, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
       Reply  Thu Mar 14 17:31:57 2013, Manasa, Update, Green Locking, X arm green locked in TEM00 
          Reply  Fri Mar 15 01:57:02 2013, Jenne, Update, Green Locking, X arm green locked in TEM00 
       Reply  Fri Mar 15 02:19:01 2013, Jenne, Update, Green Locking, Yarm ALS laser is funny / dying 
          Reply  Fri Mar 15 15:26:13 2013, Koji, Update, Green Locking, Yarm ALS laser is funny / dying 
             Reply  Mon Mar 18 14:58:15 2013, Steve, Update, Green Locking, repair or replace dead NPRO 
                Reply  Thu Mar 28 07:58:15 2013, Steve, Update, Green Locking, NPRO repair 
Message ID: 8294     Entry time: Thu Mar 14 16:41:48 2013     In reply to: 7002     Reply to this: 8296   8300
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Yarm ALS laser is funny / dying 

Quote (elog 7002, 23July2012):

Jamie and I were doing some locking, and we found that the Yarm green wasn't locking.  It would flash, but not really stay locked for more than a few seconds, and sometimes the green light would totally disappear.  If the end shutter is open, you can always see some green light on the arm transmission cameras.  So if the shutter is open but there is nothing on the camera, that means something is wrong.

I went down to the end, and indeed, sometimes the green light completely disappears from the end table.  At those times, the LED on the front of the laser goes off, then it comes back on, and the green light is back.  This also corresponds to the POWER display on the lcd on the laser driver going to ~0 (usually it reads ~680mW, but then it goes to ~40mW).  The laser stays off for 1-2 seconds, then comes back and stays on for 1-2 minutes, before turning off for a few seconds again.

Koji suggested turning the laser off for an hour or so to see if letting it cool down helps (I just turned it off ~10min ago), otherwise we may have to ship it somewhere for repairs :( 

 This is happening again to the Yend laser.  It's been fine for the afternoon, and I've been playing with the temperature.  First I have been making big sweeps, to figure out what offset values do to the actual temperature, and more recently was starting to do a finer sweep.  Using the 'max hold' function on the 8591, I have seen the beat appear during my big sweeps.  Currently, the laser temperature measurement is at the Yend, and the RF analyzer is here in the control room, so I don't know what temp it was at when the peaks appeared.

Anyhow, while trying to reaquire lock of the TEM00 mode after changing the temperature, I find that it is very difficult (the green seems misaligned in pitch), and every minute or so the light disappears, and I can no longer see the straight-through beam on the camera.  I went down to the end, and the same symptoms of LED on the laser head turning off, power out display goes to ~40mW, are happening.  I have turned off the laser as was the solution last time, in hopes that that will fix things.

Manasa has done some work to get the Xgreen aligned, so I'll switch to trying to find that beatnote for now.

ELOG V3.1.3-